i am having trouble with solving rational inequalities. especially ones like this: x+1/3 - x+1/x+2 = 0
Answer by rfer(16322)

x and -x cancel
1/3+1/x+2=0
1/x=-2 1/3
1/x=-7/3
-7x=3
x=-3/7